Shift report 20030301 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Mar 01 2003 - 16:11:21 EST

  • Next message: Apache: "Shift report 20030302 16:00-24:00"
    ------------------------------------------------------------
    Date: Sat Mar  1 16:11:20 2003, Shift: 08:00-16:00
    Supervisor: FR, with: HB
    Log book#14 pages 41 - 42
    
    Shift Summary:
    ------------------------------------------------------------
    1) Nice and long store from 10:18, still running.
    Initial rates: ZDC~10.5K, TOFW~120K, ZDC/TOFW~11.
    Rates at the end of the shift: ZDC~4K, TOFW~20K, ZDC/TOFW~5.
    
    2) 5 Data runs, all taken without T3.
    - Run 8154: FS at 8deg, 1/15A - MRS at 40deg, 1050B
    120K  trig6 events.
    - Run 8155: FS at 8deg, 1/15A - MRS at 40deg, 1050B
     246K trig6 events.
    - Run 8156: FS at 8deg, 1/4A - MRS at 40deg, 1050B
     74K trig6 events.
    - Run 8157: FS at 8deg, 1/8B - MRS at 40deg, 1050B
     300K trig7 events.
    - Run 8158: FS at 8deg, 1/8B - MRS at 40deg, 1050B,still
    running.
    
    3) There was problem with T3 tripping. We decided with to turn it off. Mainframe L1458B was off for the last 2-3
    minutes of Run8154.
    
    Plan for the next shift:
       Keep T3 off.
     - Finish current setting, Goal is 300K trig.6 events.
     - Take 200K events with FS 1/4B-field.
    
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Mar 01 2003 - 16:12:51 EST